# Restricted Spectral Gap Decomposition for Simulated Tempering Targeting Mixture Distributions

### v1

@@ Abstract.
Simulated tempering is a widely used strategy for sampling from multimodal distributions. In this paper, we consider simulated tempering combined with an arbitrary local Markov chain Monte Carlo sampler and present a new decomposition theorem that provides a lower bound on the restricted spectral gap of the algorithm for sampling from mixture distributions. By working with the restricted spectral gap, the applicability of our results is extended to broader settings such as when the usual spectral gap is difficult to bound or becomes degenerate. We demonstrate the application of our theoretical results by analyzing simulated tempering combined with random walk Metropolis–Hastings for sampling from mixtures of Gaussian distributions. We show that in fixed-dimensional settings, the algorithm’s complexity scales polynomially with the separation between modes and logarithmically with $1/\varepsilon$, where $\varepsilon$ is the target accuracy in total variation distance.
@@ 1. Introduction
For distributions that deviate significantly from log-concavity and exhibit numerous deep modes, additional techniques are often required to ensure efficient sampling. A comprehensive discussion of the fundamental challenges in sampling from multimodal distributions, as well as an overview of major types of Markov chain Monte Carlo (MCMC) algorithms designed for this purpose—including parallel tempering, mode jumping, and Wang–Landau methods—can be found in Łatuszyński et al. (2025). The recent work of Koehler et al. (2024) addresses the problem of sampling from a multimodal distribution by initializing the sampler using a small number of stationary samples. They show that, under the assumption of a $k$-th order spectral gap, the chain efficiently yields an $\varepsilon$-accurate sample in total variation (TV) distance with $\widetilde{O}(k/\varepsilon^{2})$ samples, and they further extend the result to scenarios where the score (i.e., the drift term of LMC) is estimated. Lee and Santana-Gijzen (2024) studied the sequential Monte Carlo (SMC) method (Liu and Chen, 1998) and derived the complexity result for mixture target distributions. Additionally, denoising-diffusion-based samplers have been proposed for sampling from non-log-concave targets without relying on isoperimetric inequalities (Huang et al., 2023, 2024; He et al., 2024b). These methods reverse the Ornstein–Uhlenbeck process and require estimating score functions (i.e., gradients of the log-density) via importance sampling, which becomes particularly challenging in high-dimensional settings.
@@ 2.1. Notation and Definitions
We will refer to $\mathcal{E}_{\Omega^{0}}(g,g;\Pi,K)$ as the $\Omega^{0}$-restricted Dirichlet form and omit $\Pi,K$ when they are clear from the context. When $\Omega^{0}=\Omega$, $\mathrm{SpecGap}_{\Omega^{0}}(K)$ is known as the spectral gap of $K$, and we simply write $\mathrm{SpecGap}(K)=\mathrm{SpecGap}_{\Omega}(K),\mathcal{E}(g,g)=\mathcal{E}_{\Omega}(g,g)$ and $\mathrm{Var}_{\Pi}(g)=\mathrm{Var}_{\Pi,\Omega_{0}}(g)$. Note that $\mathrm{Var}_{\Pi}(g)$ equals the variance of $g(\omega)$ with $\omega\sim\Pi$.
Simulated tempering can be viewed as a bivariate Markov chain defined as follows.
@@
Let $L\geq 2$ be an integer and $[L]=\{1,2,\dots,L\}$. Let $\lambda\in(0,1)$ and $(r_{i})_{i\in[L]}$ be constants such that $r_{i}>0$ for each $i$, and $\sum_{i=1}^{L}r_{i}=1.$ Let $M_{i}$, for each $i\in[L]$, be the transition kernel of a Markov chain with state space $\mathcal{X}$ and stationary density $p_{i}$. The simulated tempering Markov chain has state space $[L]\times\mathcal{X}$. Denote its transition kernel by
@@
$m((i,x),(i^{\prime},x^{\prime}))=\left\{\begin{array}[]{cc}(1-\lambda)m_{i}(x,x^{\prime}),&\text{ if }i=i^{\prime},x\neq x^{\prime},\\[5.0pt] \frac{\lambda}{2}a((i,x),(i^{\prime},x)),&\text{ if }i^{\prime}=i\pm 1,x=x^{\prime},\\ \end{array}\right.$
@@ 2.2. Decomposition of the Simulated Tempering Markov chain
Consider the simulated tempering Markov chain given in Definition 2. Let $\mathcal{X}^{0}\subset\mathcal{X}$ be a measurable subset of $\mathcal{X}$, and our goal is to derive a decomposition theorem for the spectral gap of $M$ restricted to $[L]\times\mathcal{X}^{0}$. We assume that the distributions $(P_{i})_{i\in[L]}$ satisfies the following assumption.
@@
Under Assumption 1, we can introduce a latent variable $J$ so that given $I=i,J=j$, $X$ is drawn from the component distribution $P_{(i,j)}$. The density of the target distribution is augmented to $p(i,j,x)=r_{i}w_{(i,j)}p_{(i,j)}(x).$ The overall strategy of our decomposition is similar to Theorem 6.3 of Ge et al. (2018) in that we directly decompose the Dirichlet form of $M$ into the Dirichlet forms of Markov chains $(M_{(i,j)})_{i\in[L],j\in[n]}$, where $M_{(i,j)}$ has stationary distribution $P_{(i,j)}$. We assume that these chains are chosen such that the following assumption is satisfied.
@@
[…]
Define the projected chain with transition matrix $\overline{M}$ on $[L]\times[n]$ by
@@
(2.8) $C_{M}=\max\left\{3\theta C_{3},\;\frac{\theta C_{1}C_{2}}{2\phi(1-\lambda)}\left({(4-\lambda)C_{3}}+2\right)\right\}.$
@@
When $\mathcal{X}^{0}=\mathcal{X}$, we obtain a bound on the spectral gap of $M$, which can be compared to the continuous-time decomposition theorem of Ge et al. (2018).
@@ 2.3. Mixing Time Bounds
3. Analysis of the Simulated Tempering Metropolis–-Hastings Algorithm
@@
where $10<\beta_{1}<\cdots<\beta_{L}=$ 1 is a sequence of inverse temperatures. The transition kernel $M^{*}_{i}$ is that of the Metropolis–Hastings algorithm targeting $p^{*}_{i}$ with a symmetric Gaussian proposal density $q(x,y)=\mathcal{N}(y;x,\eta I)$, where $\eta>0$ is the step size. The constants $r_{i}$ are set proportional to $Z_{i}/\widehat{Z}_{i}$, where $Z_{i}$ is the (unknown) normalizing constant of $p^{*}_{i}$, and $\widehat{Z}_{i}$ is its estimate. Specifically, we set
@@
For STMH, the acceptance probability in Equation (2.2) takes the form
[…]
Since it depends only on the ratio $\widehat{Z}_{i}/\widehat{Z}_{i^{\prime}}$, it is sufficient to estimate the normalizing constants (also known as partition functions, where $Z$ is viewed as a function of the temperature index $i$) up to a common multiplicative factor. We set $\widehat{Z}_{1}=1$ and estimate the other normalizing constants using the inductive strategy considered by Ge et al. (2018): for $\ell=1,\dots,L-1$, we run STMH using $\ell$ inverse temperatures $\beta_{1}<\cdots<\beta_{\ell}$ and use the samples at temperature level $\ell$ to compute the estimate $\widehat{Z}_{\ell+1}$. This estimation procedure is summarized in Algorithm 2 in Appendix A. After obtaining all estimates $(\widehat{Z}_{i})_{i\in[L]}$, we run STMH using all $L$ temperature levels, and the samples collected at the $L$-th temperature level can be used to approximate the original target density $p^{*}(x)\propto e^{-f(x)}$. Algorithm 1 in Appendix A summarizes the STMH algorithm given in Definition 4, assuming all partition functions are known.
@@
The time complexity in Theorem 5 is polynomial in $D$ and logarithmic in the inverse accuracy $1/\varepsilon$. In fixed-dimensional settings—where $d$ is constant—this result provides high-accuracy guarantees, showing that STMH is particularly effective for sampling from such mixture distributions. For comparison, the proximal sampler incurs a time complexity with exponential dependence on $D$, since the log-Sobolev constant and Poincaré constant for a Gaussian mixture distribution decays exponentially in $D$; see Schlichting (2019). To further illustrate the scaling behavior of the proximal sampler, consider a mixture of two Gaussians with identical covariance matrices and means located at $(-m/2,\dots,-m/2)$ and $(m/2,\dots,m/2)$, respectively. In this case, the separation becomes $D=m\sqrt{d}$, which implies that the complexity of the proximal sampler scales exponentially in both $m$ and $d$. The simulated tempering Langevin Monte Carlo (STLMC) algorithm of Ge et al. (2018) scales polynomially with $D$ and $d$ but also has a polynomial dependence on $1/\varepsilon$, while STMH achieves a more favorable logarithmic dependence on $\varepsilon$. Table 1 summarizes the theoretical complexity of STMH in comparison with several other sampling algorithms. It can be seen that no algorithm dominates STMH in terms of the complexity dependence on $D$ and $1/\varepsilon$. In the proof of Theorem 5, it is shown that the time complexity of STMH has exponential dependence on $d$, which is likely due to the use of the random walk Metropolis–Hastings sampler for each local chain $M_{i}$. We conjecture that by replacing it with proximal sampler (Chen et al., 2022; He et al., 2024a) or Metropolis-adjusted Langevin algorithm (Wu et al., 2022), the resulting simulated tempering algorithm may achieve a better complexity dependence on $d$.
@@ 4. Simulation Study
where $I_{2}$ is the $2\times 2$ identity matrix and ${1_{2}}=(1,1)^{\top}$. The parameter $D$ controls the separation between the two components. We vary the parameter $D$ to explore how increasing the separation between the modes affects the convergence behavior of the algorithm. For each value of $D$, we run the STMH algorithm with the parameters specified in Appendix C.4.1 and initialized at $(10,10)$. To assess convergence, we monitor how quickly the empirical mean of the samples, denoted by $\hat{\mu}$, approaches the true mean of the target distribution, $(0,0)$. If the chain has not yet mixed, it tends to spend more time near one mode, resulting in an empirical mean that deviates from the target mean. Moreover, according to  Nishiyama (2022), when the empirical and target distribution have the same covariance matrix, we can lower bound their TV distance by $\|\hat{\mu}\|^{2}/(C+\|\hat{\mu}\|^{2})$ for some constant $C>0$, which is a monotone increasing function of $\|\hat{\mu}\|^{2}$. This justifies the use of $\|\hat{\mu}\|$ as a measure of convergence. To reduce variability, we repeat the simulation $20,000$ times for each fixed $D$ and average the empirical means over all runs. All simulations were performed on a standard consumer-grade CPU with parallelization and completed within approximately three hours.
In the left panel of Figure 1, we plot the number of steps $N$ required for $\|\hat{\mu}\|$ to drop below 0.1. It can be seen that this number increases approximately linearly with $D^{2}$, consistent with the theoretical bound in Theorem 5. To analyze how the mixing time depends on the threshold $\varepsilon$, we fix $D=30$ and plot how $\|\hat{\mu}\|$ varies with the number of steps $N$ in the right panel of Figure 1. We see that $N$ and $\log^{2}(1/\|\hat{\mu}\|)$ have an approximately linear relationship, which is again consistent with the theoretical bound in Theorem 5. [Uncaptioned image] [Uncaptioned image]
Figure 1. Left: number of steps until $\|\widehat{\mu}\|<0.1$ versus $D^{2}$. Right: $\log^{2}\left(1/\|\widehat{\mu}\|\right)$ versus the number of steps $N$ for $D=30$.

### v2

@@ Abstract.
Simulated tempering is a widely used strategy for sampling from multimodal distributions. In this paper, we consider simulated tempering combined with an arbitrary local Markov chain Monte Carlo sampler and present a new decomposition theorem that provides a lower bound on the restricted spectral gap of the algorithm for sampling from mixture distributions. By working with the restricted spectral gap, the applicability of our results is extended to broader settings such as when the usual spectral gap is difficult to bound or becomes degenerate. We demonstrate the application of our theoretical results by analyzing simulated tempering combined with random walk Metropolis–Hastings for sampling from mixtures of Gaussian distributions. Our complexity bound scales polynomially with the separation between modes, logarithmically with $1/\varepsilon$, where $\varepsilon$ denotes the target accuracy in total variation distance, and exponentially with the dimension $d$.
@@ 1. Introduction
For distributions that deviate significantly from log-concavity and exhibit numerous deep modes, additional techniques are often required to ensure efficient sampling. A comprehensive discussion of the fundamental challenges in sampling from multimodal distributions, as well as an overview of major types of Markov chain Monte Carlo (MCMC) algorithms designed for this purpose—including parallel tempering, mode jumping, and Wang–Landau methods—can be found in Łatuszyński et al. (2025). The recent work of Koehler et al. (2024) addresses the problem of sampling from a multimodal distribution by initializing the sampler using a small number of stationary samples. They show that, under the assumption of a $k$-th order spectral gap, the chain efficiently yields an $\varepsilon$-accurate sample in total variation (TV) distance with $\widetilde{O}(k/\varepsilon^{2})$ samples, and they further extend the result to scenarios where the score (i.e., the drift term of LMC) is estimated. Lee and Santana-Gijzen (2024) studied the sequential Monte Carlo (SMC) method (Liu and Chen, 1998; Del Moral et al., 2006; Chopin et al., 2020; Syed et al., 2024) and derived the complexity result for mixture target distributions. Additionally, denoising-diffusion-based samplers have been proposed for sampling from non-log-concave targets without relying on isoperimetric inequalities (Huang et al., 2023, 2024; He et al., 2024b). These methods reverse the Ornstein–Uhlenbeck process and require estimating score functions (i.e., gradients of the log-density) via importance sampling, which becomes particularly challenging in high-dimensional settings.
@@ 2.1. Notation and Definitions
We will refer to $\mathcal{E}_{\Omega^{0}}(g,g;\Pi,K)$ as the $\Omega^{0}$-restricted Dirichlet form and omit $\Pi,K$ when they are clear from the context. When $\Omega^{0}=\Omega$, $\mathrm{SpecGap}_{\Omega^{0}}(K)$ is known as the spectral gap of $K$, and we simply write $\mathrm{SpecGap}(K)=\mathrm{SpecGap}_{\Omega}(K),\mathcal{E}(g,g)=\mathcal{E}_{\Omega}(g,g)$ and $\mathrm{Var}_{\Pi}(g)=\mathrm{Var}_{\Pi,\Omega_{0}}(g)$. Note that $\mathrm{Var}_{\Pi}(g)$ equals the variance of $g(\omega)$ with $\omega\sim\Pi$. Intuitively, the spectral gap quantifies how rapidly a Markov chain mixes: a larger gap corresponds to faster convergence to stationarity distribution. The restricted spectral gap generalizes this idea by measuring the rate of mixing in a subset $\Omega^{0}$ of the state space.
Simulated tempering can be viewed as a bivariate Markov chain. The key idea is to augment the state space with a temperature index that allows the chain to move between a family of distributions, ranging from the target distribution (low temperature) to flattened versions (high temperature). At higher temperatures the landscape is smoother, enabling easier transitions between modes, while samples from the target are obtained by collecting states only at the lowest temperature. Formally, it is defined as follows.
@@
Let $L\geq 2$ be an integer and $L]=\{1,2,\dots,L\}$. Let $\lambda\in(0,1)$ and $(r_{i})_{i\in[L]}$ be constants such that $r_{i}>0$ for each $i$, and $\sum_{i=1}^{L}r_{i}=1.$ Let $M_{i}$, for each $i\in[L]$, be the transition kernel of a Markov chain with state space $\mathcal{X}$ and stationary density $p_{i}$. The simulated tempering Markov chain has state space $L]\times\mathcal{X}$. Denote its transition kernel by
@@
[…]
@@ 2.2. Decomposition of the Simulated Tempering Markov chain
Consider the simulated tempering Markov chain given in Definition 2. Let $\mathcal{X}^{0}\subset\mathcal{X}$ be a measurable subset of $\mathcal{X}$, and our goal is to derive a decomposition theorem for the spectral gap of $M$ restricted to $L]\times\mathcal{X}^{0}$. We assume that the distributions $(P_{i})_{i\in[L]}$ satisfies the following assumption.
@@
For simplicity, in Assumption 1 we assume that each stationary density $p_{i}$ can be expressed as a mixture of simpler component densities, an assumption also made in Ge et al. (2018). Under Assumption 1, we can introduce a latent variable $J$ so that given $I=i,J=j$, $X$ is drawn from the component distribution $P_{(i,j)}$. The density of the target distribution is augmented to $p(i,j,x)=r_{i}w_{(i,j)}p_{(i,j)}(x).$
The overall strategy of our decomposition is similar to Theorem 6.3 of Ge et al. (2018) in that we directly decompose the Dirichlet form of $M$ into the Dirichlet forms of Markov chains $(M_{(i,j)})_{i\in[L],j\in[n]}$, where $M_{(i,j)}$ has stationary distribution $P_{(i,j)}$. We assume that these chains are chosen such that the following assumption is satisfied.
@@
Assumption 2 assumes a lower bound on the mixing rate of the local chains $M_{(i,j)}$, each targeting a single component $p_{(i,j)}$. Efficient mixing within each component is essential, as poor local mixing can limit global exploration and slow the overall convergence of the simulated tempering chain.
@@
$a((i,j,x),(i^{\prime},j^{\prime},x))=\min\left\{\frac{r_{i^{\prime}}w_{(i^{\prime},j^{\prime})}p_{(i^{\prime},j^{\prime})}(x)}{r_{i}w_{(i,j)}p_{(i,j)}(x)},\;1\right\},\quad p_{i}(j|x)=\frac{w_{(i,j)}p_{(i,j)}(x)}{p_{i}(x)}.$
Define the projected chain with transition matrix $\overline{M}$ on $L]\times[n]$ by
@@
Assumption 3 assumes a lower bound on the mixing rate of the projected chain. Intuitively, if the components are well-separated, transitions between them becomes rare, resulting in slow mixing of the projected chain and, consequently, slow overall convergence of the simulated tempering chain. Conversely, when components are closer and transitions are more likely, the chain mixes more rapidly.
@@
(2.8) $C_{M}=\max\left\{3\theta C_{3},\;\frac{\theta C_{1}C_{2}}{\phi(1-\lambda)}\left({(2+\lambda)C_{3}}+1\right)\right\}{\mathcal{E}_{[L]\times\mathcal{X}^{0}}}(g,g)$}.
@@
When $\mathcal{X}^{0}=\mathcal{X}$, we obtain a bound on the spectral gap of $M,$ which is similar in spirit to the continuous-time decomposition theorem of Ge et al. (2018), though the two results are not directly comparable due the discrete-time setting and our construction of $\overline{M}$.
@@ 2.3. Mixing Time Bounds
The constant $C_{M}$ controls the rate of convergence of the simulated tempering chain to its stationary distribution in total variation distance; larger values of $C_{M}$ correspond to slower convergence.
3. Analysis of the Simulated Tempering Metropolis–-Hastings Algorithm for Multivariate Gaussian Mixtures
@@
where $0<\beta_{1}<\cdots<\beta_{L}=1$ is a sequence of inverse temperatures. The transition kernel $M^{*}_{i}$ is that of the Metropolis–Hastings algorithm targeting $p^{*}_{i}$ with a symmetric Gaussian proposal density $q(x,y)=\mathcal{N}(y;x,\eta I)$, where $\eta>0$ is the step size. The constants $r_{i}$ are set proportional to $Z_{i}/\widehat{Z}_{i}$, where $Z_{i}$ is the (unknown) normalizing constant of $p^{*}_{i}$, and $\widehat{Z}_{i}$ is its estimate. Specifically, we set
@@
We define $r_{i}\propto Z_{i}/\widehat{Z}_{i}$ because the true normalizing constants $Z_{i}$ (also known as partition functions, where $Z$ is viewed as a function of the temperature index $i$) are typically unknown in practice. When implementing the STMH algorithm, the acceptance probability in Equation (2.2) is given by
$a\big((i,x),(i^{\prime},x)\big)=\min\left\{\frac{\widehat{Z}_{i}\exp(-\beta_{i^{\prime}}f(x))}{\widehat{Z}_{i^{\prime}}\exp(-\beta_{i}f(x))},\,1\right\}.$
This choice ensures that the acceptance probability depends only on the estimated normalizing constants $\widehat{Z}_{i}$ and not the true values $Z_{i}$, thereby making the algorithm implementable even when $Z_{i}$ are unknown. Since acceptance probability depends only on the ratio $\widehat{Z}_{i}/\widehat{Z}_{i^{\prime}}$, it is sufficient to estimate the normalizing constants up to a common multiplicative factor. We set $\widehat{Z}_{1}=1$ and estimate the other normalizing constants using the inductive strategy considered by Ge et al. (2018): for $\ell=1,\dots,L-1$, we run STMH using $\ell$ inverse temperatures $\beta_{1}<\cdots<\beta_{\ell}$ and use the samples at temperature level $\ell$ to compute the estimate $\widehat{Z}_{\ell+1}$. This estimation procedure is summarized in Algorithm 2 in Appendix A. After obtaining all estimates $(\widehat{Z}_{i})_{i\in[L]}$, we run STMH using all $L$ temperature levels, and the samples collected at the $L$-th temperature level can be used to approximate the original target density $p^{*}(x)\propto e^{-f(x)}$. Algorithm 1 in Appendix A summarizes the STMH algorithm given in Definition 4, assuming all partition functions are known.
@@
The time complexity in Theorem 5 is polynomial in the separation parameter $D$, logarithmic in the inverse accuracy $1/\varepsilon$, and exponential in the dimension $d$. In fixed-dimensional settings—where $d$ is constant—this result provides high-accuracy guarantees, showing that STMH is particularly effective for sampling from such mixture distributions. For comparison, the proximal sampler incurs a time complexity with exponential dependence on $D$, since the log-Sobolev constant and Poincaré constant for a Gaussian mixture distribution decays exponentially in $D$; see Schlichting (2019). To further illustrate the scaling behavior of the proximal sampler, consider a mixture of two Gaussians with identical covariance matrices and means located at $(-m/2,\dots,-m/2)$ and $(m/2,\dots,m/2)$, respectively. In this case, the separation becomes $D=m\sqrt{d}$, which implies that the complexity of the proximal sampler scales exponentially in both $m$ and $d$. The simulated tempering Langevin Monte Carlo (STLMC) algorithm of Ge et al. (2018) admits an upper bound that scales polynomially with $D$, $d$, and $1/\varepsilon$. In contrast, the upper bound we obtain for STMH exhibits only a logarithmic dependence on $1/\varepsilon$, representing a state-of-the-art theoretical guarantee among existing known upper bounds. Table 1 summarizes the theoretical complexity of STMH in comparison with several other sampling algorithms. It can be seen that no algorithm dominates STMH in terms of the complexity dependence on $D$ and $1/\varepsilon$. Our upper bound on the time complexity of STMH has exponential dependence on $d$, which is likely due to the use of the random walk Metropolis–Hastings (RWMH) sampler for each local chain $M_{i}$. We conjecture that by replacing it with proximal sampler (Chen et al., 2022; He et al., 2024a) or Metropolis-adjusted Langevin algorithm (Wu et al., 2022), the resulting simulated tempering algorithm may achieve a better complexity dependence on $d$.
Remark 6.
Theorem 5 establishes convergence guarantees for STMH when sampling from a mixture of Gaussians with a shared covariance matrix. This result can be naturally extended to target distributions that are sufficiently close to such mixtures, following an approach similar to that of Ge et al. (2018). In such cases, the time complexity would additionally depend on closeness between the actual distribution and the Gaussian mixture approximation. We also anticipate that similar techniques can be adapted to handle mixtures of log-concave distributions that are translates to each other, or more broadly, distributions that are well-approximated by such mixtures. However, as demonstrated by Ge et al. (2018), some seemingly mild violations of the assumptions, such as component covariance matrices differing by a constant factor, can lead to exponential time complexity for any reasonable algorithm with similar guarantees.
@@ 4. Simulation Study
where $I_{2}$ is the $2\times 2$ identity matrix and ${1_{2}}=(1,1)^{\top}$. The parameter $D$ controls the separation between the two components. We vary the parameter $D$ to explore how increasing the separation between the modes affects the convergence behavior of the algorithm. For each value of $D$, we run the STMH algorithm with the parameters specified in Appendix C.4.1 and initialized at $(10,10)$. To assess convergence, we monitor how quickly the empirical mean of the samples, denoted by $\hat{\mu}$, approaches the true mean of the target distribution, $(0,0)$. If the chain has not yet mixed, it tends to spend more time near one mode, resulting in an empirical mean that deviates from the target mean. Moreover, according to  Nishiyama (2022), when the empirical and target distribution have the same covariance matrix, we can lower bound their TV distance by $\|\hat{\mu}\|^{2}/(C+\|\hat{\mu}\|^{2})$ for some constant $C>0$, which is a monotone increasing function of $\|\hat{\mu}\|^{2}$. This justifies the use of $\|\hat{\mu}\|$ as a measure of convergence. To reduce variability, we repeat the simulation $20,000$ times for each fixed $D$ and average the empirical means over all runs. To provide a benchmark, we also run the baseline Metropolis–Hastings (MH) algorithm under the same setup and compare its convergence behavior with that of STMH. All simulations were performed on a standard consumer-grade CPU with parallelization and completed within approximately six hours.
[…]